OpenStudy (anonymous):

A wooden artifact from an ancient tomb contains 30 percent of the carbon-14 that is present in living trees. How long ago was the artifact made? (The half-life of carbon-14 is 5730 years.) Your answer is ? years.

OpenStudy (anonymous):

formula is\[Q=Q_0(\frac{1}{2})^{\frac{t}{5700}}\]

OpenStudy (anonymous):

so you need to solve \[.30=(\frac{1}{2})^{\frac{t}{5700}}\] for t

OpenStudy (anonymous):

take the log \[ln(.3)=\frac{t}{5700}ln(\frac{1}{2})\] solve for t to get \[t=5700\frac{ln(.3)}{ln(.5)}\]

OpenStudy (anonymous):

now use a calculator

OpenStudy (anonymous):

i get 9900